The average of x,y, andz is 80. If two of the numbers are 74 and 78, then what is the other number.
Mrs Jones is installing wall to wall carpet in a room that measures 12 1/2ft by 9 2/3 ft. how much will it cost if the carpet is $26.00 per square yard?

Thank you

Answer by stanbon(48568)

The average of x,y, andz is 80. If two of the numbers are 74 and 78, then what is the other number.
---
EQUATION:
(74 + 78 + z)/3 = 80
152 + z = 240
z = 88
============
Mrs Jones is installing wall to wall carpet in a room that measures 12 1/2ft by 9 2/3 ft. how much will it cost if the carpet is $26.00 per square yard?
Area = [12 1/2][9 2/3] = 725/6 square feet
---------------------------
Cost = (725/6)*26 = $3141.67
==============================
Cheers,
Stan H.